OpenStudy (anonymous):

What is the radical expression of 4 OVER a^9

OpenStudy (anonymous):

\[\frac{4}{ a^9} = \sqrt{\frac{4^2}{ (a^9)^2}}= \sqrt{\frac{16}{ a^{18}}}\]

OpenStudy (anonymous):

@troyandthor

OpenStudy (anonymous):

The choices are: a. \[4a^9\] b. \[9a^4\] c. \[4\sqrt{a^9}\] b. \[9\sqrt{a^}\]

OpenStudy (anonymous):

b. \[9\sqrt{a^4}\]

OpenStudy (anonymous):

@dpasingh